# COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Gravitational Lenses in DECaLS

### 17 Jan 2024 v2

@@ 3.2 Stellar Luminosity and Mass
We conducted SED modelling using the photometry obtained in Section 2.4 with Prospector (Johnson et al., 2021; Conroy et al., 2009; Conroy & Gunn, 2010; Foreman-Mackey et al., 2013). We fit a parametric SED with three assumed parameters and five free parameters. We assumed a Chabrier initial mass function (IMF) (Chabrier, 2003), Calzetti attenuation curve for dust (Calzetti et al., 2000), and delayed-tau star-formation history (SFH) model (Carnall et al., 2019). The five free parameters that were optimized are summarized in Table 1. We utilized emcee as implemented in Prospector with 84 walkers for a total of 6720 iterations, taking only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) as representing the posterior distribution for each of the variables of interest in order to eliminate the significant burn-in sequence of the fit. After fitting, we adopted the 50th percentile of the posterior distribution as the value for each of the five parameters, and the greater difference between the 84th-50th percentile and 50th-16th percentile values as the error. Observed-frame spectra for the systems shown in Figure 1 after fitting are shown in Figure 2.
To derive luminosity, we drew random sample of 1024 chains from the obtained posterior distribution and generated an SED for each. We integrated over the rest-frame wavelength range of 3000Å-7000Å, since this corresponds to the longest wavelength range sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts). After integrating a given SED to obtain flux, we converted the observed flux to a luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random draws as the luminosity and its and error respectively for each system.
@@ 4 Results
As Figure 3 shows, the BCG redshift, enclosed total mass, and enclosed luminosity are correlated. Indeed, we note that the most massive clusters in this work contain a large amount of stars and thus also have a large luminosity. Using emcee, we fit a plane-of-best-fit using likelihood maximization taking into account the uncertainty in variables on all three axes to describe this correlation where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$, yielding:
@@
This simple and analyic relationship derived from a sample of known strong lenses, considering only quantities derived from observations, may have utility in differentiating bona fide strong lenses from non-lenses in large samples of candidate lenses where the evidence of lensing is not nearly as robust as the sample of clusters analyzed here. Moreover, this correlation offers a future opportunity to test simulations against observations, particularly those that aim to generate realistic galaxy populations in dense cluster environments.
Figure 3: A 3D visualization of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles with a plane of best fit shown in black as well. Uncertainties are represented as red, blue, and green lines respectively.
@@ 5.1 Scaling Relations
Besides the enclosed total mass and luminosity, we can also infer the current enclosed stellar mass at the redshift of observation for each system using the posterior distribution of each SED fit generated with Prospector. This was obtained by correcting the derived enclosed stellar mass in each system by the surviving mass fraction returned by Prospector, since the amount of stellar mass that we see at the time of observation is less than the total stellar mass formed over the lifetime of the galaxy (e.g., Li et al. 2017). While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
@@ 5.1.1 Total Mass — Stellar Mass
Figure 4: The total mass contained within the Einstein aperture for each system plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction. Points in panels I and III are color-coded according to the luminosity of the system in units of solar luminosities from stellar light within the Einstein aperture as measured in the rest-frame wavelength range from 3000Å to 7000Å. Points in panels II and IV only represent those systems in which multiple cluster members fall within the Einstein aperture. Linear regression lines in all four subpanels are drawn from the posterior distribution of their respective MCMC.
In Panel I of Figure 4, we see a clearly positive correlation between the enclosed stellar mass and the enclosed total mass for each system. By performing a linear regression with emcee using a likelihood maximization that accounts for uncertainties in variables on both axes, we find that the slope of this correlation is $0.443\pm 0.035$ from the posterior distribution of the MCMC fit. We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall systematic scatter. In Table 2 of Kravtsov et al. (2018), they found the stellar mass of solely the BCG and $M_{500}$ (White, 2001) to be correlated with a logarithmic slope of $0.33\pm 0.11$. In addition, Kravtsov et al. (2018) found the stellar mass of the entire cluster and $M_{500}$ to be correlated with a logarithmic slope of $0.59\pm 0.08$. The fact that the slope for this work falls between the two measurements quoted by Kravtsov et al. (2018) is consistent with how our methodology samples a physical regime between exclusively the BCG and the entire cluster.
@@ 5.1.2 Total Mass — Stellar-to-Total Mass Fraction
In Panel III of Figure 4, we compare the enclosed stellar-to-total mass fraction against the enclosed total mass and see a clearly negative correlation (again with the same colormapping as Panel I to emphasize intrinsic scatter). We find that the slope of this correlation is $-0.563\pm 0.035$. Andreon (2010) measured 52 clusters over a slightly larger mass range and found that the stellar mass fraction depended on halo mass with a slope of $-0.55\pm 0.08$ when measuring to a physical extent of $R_{200}$. Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.561\pm 0.050$. We again find that the multiple-galaxy systems are statistically indistinguishable from the slope for the entire sample shown in Panel III and in Andreon (2010). While the linear fit derived by Andreon (2010) has a slightly larger normalization, this is to be expected because the cluster masses are constrained out to $R_{200}$.
@@ 5.2 Photometric Bias and Error
Deriving light via directly integrating over the Einstein aperture presents some challenges. Since we identify red-sequence cluster members by using observed color as a proxy for cluster membership, it is possible that we may have included intrinsically-bluer galaxies at a higher redshift that appear to be the same color as the cluster members. Conversely, we may have excluded intrinsically-bluer bona fide cluster members. However, star-forming cluster members have less stellar mass and are intrinsically rarer in cluster cores (e.g., Dressler et al. 1997, 2004), and we do not expect that these missed cluster galaxies introduce a significant loss to the measured cluster light. In addition, bluer galaxies at a higher redshift appear intrinsically fainter, which are also unlikely to bias our methodology of measuring cluster light. Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying galaxy flux is typically minimal due to the limited angular size of such stars.
@@
In general, we expect that all of the unmasked flux from cluster cores at the angular scales considered in this work originates from genuine cluster members. Any modest background correction that could be made will also be effectively captured in the redshift dependence of the fits (e.g., Figure 3). While we must also consider associated galaxy populations in the immediate large-scale structure when considering galaxy clusters projected onto the sky—as such structures may also be projected onto the cluster core—those galaxies will also contribute to the lensing effect and will have their mass and light effectively captured by the methodology presented here.
@@
There are also extremely few examples of systems whose Einstein radius is less than $\sim 2^{\prime\prime}$. To some extent, this is a consequence of the chosen sample of galaxy clusters and groups, which we would generally expect to have larger Einstein radii. Systems whose arcs are separated from the central galaxy at small radii may be located directly within the discernable BCG light. This would not only make them hard to visually notice, but it would also make them unsuitable for this analysis because the source light would be significantly blended together with the lensing cluster light. However, we caution that there is a large scatter between the total halo mass and Einstein radius (Fox et al., 2022). While we thus expect this work to undersample the small-Einstein-radius regime, it is not apparent that this significantly alters any of our results.
@@
In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $grz$ bands to $\sim$ 24.7, 23.9, and 23.0 respectively in the average case of two exposures per filter (Dey et al., 2019). Clusters higher than redshift $\sim 1$ are barely visible even in $\mathit{z}$-band filter images in the current data and are simply a rarer type of lensing system in the universe (Li et al., 2019). Both the data depth and the intrinsic redshift distribution of lenses shape the redshift distribution in this work.
@@ 5.3 Alternative Mass Measurement
In constraining mass, we also considered a second method for measuring the cluster-centric mass elucidated in Remolina González et al. (2021a) using the parametric lens modelling software LENSTOOL (Jullo et al., 2007) to measure the enclosed mass as opposed to the simple evaluation of Equations (1), (2), and (3). For the first 35 clusters that were analyzed, we generated single-halo lens models by using the three coordinates corresponding to the locations of the brightest clumps in each lensed arc as positional constraints with a single pseudo-isothermal elliptical mass distribution (PIEMD, Kassiola & Kovner 1993) locked to the center of the BCG. Mass estimates were then obtained from the resulting best-fit models according to the methodology described in Remolina González et al. (2021a).
Figure 5: A comparison of the single-halo LENSTOOL generated mass estimates ($M_{SHM}$) and the parametrically generated mass estimates ($M_{\Sigma}$) for the first 35 clusters analyzed in this work. The black line represents an ideal 1-1 correspondence, and the blue lines represent the uncertainty on the real correspondence.
As expected from Remolina González et al. (2021b), the two different methodologies show similar results, and the two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ as shown in Figure 5. We abandoned the LENSTOOL approach throughout the rest of the work, i.e. we find no evidence for a significant difference between these two methods.
In theory, lens modelling has the potential to be the most accurate mass-constraining methodology. However, this requires high-spatial-resolution imaging data and ample time. The process of iteratively refining just one lens model may take upwards of an hour for even a single relatively-simple system, which is untenable when the number of candidate systems reaches into the thousands. Moreover, the accuracy of measuring mass with lens modelling is sensitive to the choice of defining multiple image families along the lensed arc in a way that simple circle fitting is not. These essential positional constraints are often unresolved with even the highest-quality ground-based data. Conversely, the positional constraints that inform the Einstein radius ultimately used by Equations (1), (2), and (3) need not be extremely accurate with respect to exactly constraining image family positions. As long as the positional constraints are reasonably placed within the arcs, something which is resolved by ground-based data, it is accurate to within the systematic uncertainty.
@@ 6 Summary and Conclusions
This paper presents measurements of mass and light in 177 strong-lensing systems at cluster and group mass scales. By directly obtaining the Einstein radius and BCG redshift for each system, we compute the Einstein-radius-enclosed total mass parametrically. We use the enclosed $grz$ photometry and Prospector to parameterize the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$. We find that the enclosed total mass and stellar mass for the entire sample of strong lensing systems considered in this work are correlated with a logarithmic slope, normalization, and scatter of $0.443\pm 0.035$, $5.958\pm 0.454$, and $0.173\pm 0.013$ respectively. We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and scatter of $-0.563\pm 0.035$, $6.034\pm 0.457$, and $0.161\pm 0.013$ respectively. Enormous volumes of imaging data from new space-based and ground-based surveys at deeper and sharper photometric limits are on the horizon. Probing these upcoming datasets with both visual and machine-learning driven searches for new strong lenses at a variety of scales is inevitable. The correlations shown in this work should have utility in filtering rankings of possible strong lensing systems as well as offering another measure of cluster cores that can be compared to cosmological simulations.

### 24 Jan 2024 v3

@@ 3.2 Stellar Luminosity and Mass
We conducted parametric SED fitting using the photometry obtained in Section 2.4 with Prospector (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson et al., 2021, 2023); five free parameters and three assumed parameters parameterized each SED. A summary of all eight parameters used in fitting can be found in Table 1. Utilizing emcee (Foreman-Mackey et al., 2013) as implemented in Prospector with 84 walkers for a total of 6720 iterations, only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) were taken as representing the posterior distribution for each of the five free parameters in each system in order to eliminate the significant burn-in sequence of the fitting process. After fitting, we adopted the 50th percentile of the posterior distribution as the value for each free parameter, and the greater difference between the 84th-50th percentile and 50th-16th percentile values as the error. Observed-frame SEDs for the systems shown in Figure 1 after SED fitting are shown in Figure 2.
To derive luminosity, we generated new SEDs with a random sample of 1024 free-parameter vectors from the posterior distributions of the best-fit SED in each system. We integrated over the rest-frame wavelength interval of 3000Å to 7000Å, since this corresponds to the longest wavelength interval sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts). This range also allows us to sample the rest-frame 4000Å break while recovering flux in longer wavelengths which are significantly more luminous in cluster galaxies. After integrating the SEDs, we converted flux to luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random parameter vectors as the luminosity and its error respectively for each system.
@@ 4 Results
As Figure 3 shows, the BCG redshift, enclosed total mass, and enclosed luminosity are correlated. Indeed, we note that the most massive clusters in this work also possess a large stellar mass and thus have a large luminosity. Using emcee, we constrained a plane-of-best-fit using a likelihood maximization estimation taking into account the error in data on all three axes to describe this correlation—where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$—yielding:
@@
This simple and analyic relationship derived from a sample of known strong lenses, considering only quantities derived from observations, may have utility in differentiating bona fide strong lenses from non-lenses in large samples of candidate lenses where the evidence of lensing is not nearly as robust as the sample of clusters analyzed here. Moreover, this correlation offers a future opportunity to test simulations against observations—particularly those that aim to generate realistic galaxy populations in dense cluster environments.
Figure 3: A 3D visualization of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles. A plane-of-best-fit is also shown in black as well. Errors in data on each axis are represented as red, blue, and green lines respectively.
@@ 5.1 Scaling Relations
Besides the enclosed total mass and luminosity, we can also infer the observed enclosed stellar mass at the redshift of observation for each system using the stellar-mass posterior distribution from each SED constrained with Prospector. We obtained this by correcting the derived enclosed stellar mass formed in each system with the surviving mass fraction returned by Prospector. This is necessary because the amount of stellar mass that we see at the redshift of observation is less than the total stellar mass formed over the lifetime of the galaxy (e.g., Li et al. 2017)—the latter being what Prospector parametrically fits. While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong-lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
@@ 5.1.1 Total Mass — Stellar Mass
Figure 4: The total mass enclosed within the Einstein aperture for each system on the x-axis plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction on the y-axes. Points in Panels I and III are color coded according to the luminosity of the system in units of solar luminosities from stellar light enclosed within the Einstein aperture as measured within the rest-frame wavelength interval of 3000Å to 7000Å. Points in Panels II and IV only represent systems in which multiple cluster members fall within the Einstein aperture. Linear regression lines in all four Panels are drawn from the posterior distributions of their respective MCMC.
In Panel I of Figure 4, we see a clearly positive correlation between the enclosed total mass and the enclosed stellar mass for all systems. By performing a linear regression with emcee using a likelihood maximization estimation that accounts for errors in data on both axes, we find that the slope of this correlation is $0.443\pm 0.035$ from the posterior distribution of the MCMC fit. We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall intrinsic scatter. In Table 2 of Kravtsov et al. (2018), they found $M_{500}$ (White, 2001) and the stellar mass of solely the BCG to be correlated with a logarithmic slope of $0.33\pm 0.11$. In addition, Kravtsov et al. (2018) found $M_{500}$ and the stellar mass of the entire cluster to be correlated with a logarithmic slope of $0.59\pm 0.08$. The fact that the slope for this work falls between the two slopes quoted by Kravtsov et al. (2018) is consistent with how our methodology samples a physical regime between exclusively the BCG and that of the entire cluster.
@@ 5.1.2 Total Mass — Stellar-to-Total Mass Fraction
In Panel III of Figure 4, we plot the enclosed stellar-to-total mass fraction against the enclosed total mass and see a clearly negative correlation (again with the same colormapping as Panel I to emphasize intrinsic scatter). Also using emcee, we find that the slope of this correlation is $-0.563\pm 0.035$. Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.561\pm 0.050$. We again find that the slope for the multiple-galaxy systems are statistically indistinguishable from the slope for the entire sample shown in Panel III. Andreon (2010) measured 52 clusters over a slightly larger mass range and found that the stellar mass fraction depended on halo mass with a slope of $-0.55\pm 0.08$ when measuring out to a physical extent of $R_{200}$. Both of the slopes in Panels III and IV are statistically similar to Andreon (2010), and while the linear fit derived by Andreon (2010) has a slightly larger normalization, this is to be expected because the cluster masses are constrained out to $R_{200}$.
@@ 5.2 Photometric Bias and Error
Deriving light in Section 2.4 via directly integrating over the Einstein aperture presents some challenges. Since we identified red-sequence cluster members by using observed color as a proxy for cluster membership, it is possible that we may have included intrinsically bluer galaxies at redshift $z>z_{{}_{BCG}}$ that appear to be the same color as the cluster members. Conversely, we may have excluded intrinsically bluer bona fide cluster members. However, star-forming cluster members have less stellar mass and are intrinsically rarer in cluster cores (e.g., Dressler et al. 1997, 2004), and we do not expect that these missed cluster galaxies introduce a significant loss to the measured cluster light. In addition, bluer galaxies at redshift $z>z_{{}_{BCG}}$ appear relatively fainter, which are also unlikely to bias our methodology of measuring cluster light. Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying lens-galaxy flux is typically minimal due to the limited angular size of such stars.
@@
In general, we expect that all of the unmasked flux from cluster cores at the angular scales considered in this work originates from genuine cluster members. Any modest background correction that could be made will also be effectively captured in the redshift dependence of the fits (e.g., Figure 3). While we must also consider associated galaxy populations in the immediate large-scale structure when considering galaxy clusters projected onto the sky—as such structures may also be projected onto the cluster core—those galaxies will also contribute to the lensing effect and will have their mass and light captured by the methodology presented here.
@@
We also note that our sample contains extremely few examples of systems which have a measured Einstein radius of $\theta_{E}\lessapprox 2^{\prime\prime}$. To some extent, this is a consequence of our chosen sample of cluster-scale lenses, which we would generally expect to have larger Einstein radii. However, we caution that there exists a large scatter between the total halo mass and Einstein radius (Fox et al., 2022). Systems whose arcs are separated from the central galaxy at small radii may also be located directly within the discernable BCG light. This would not only make them hard to visually notice, but it would also make them unsuitable for this analysis because the source light would be significantly blended together with the lensing cluster light. While we thus expect this work to undersample the small-Einstein-radius regime, it is not apparent that this significantly alters our findings.
@@
In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $g$, $r$, and $z$-band filters to $\approx$ 24.7, 23.9, and 23.0 respectively in the average case of two exposures per filter (Dey et al., 2019). Systems with a lens redshift $z\gtrapprox 1$ are barely visible even in $\mathit{z}$-band LS DR9 imaging data, and they are also simply a rarer type of lensing system in the universe (Li et al., 2019). Both the data depth and intrinsic redshift distribution of lenses shape the redshift distribution in this work.
@@ 5.3 Alternative Mass Measurement
In constraining mass, we also considered a second method for measuring cluster-centric mass elucidated in Remolina González et al. (2021a) using the parametric lens-modelling software LENSTOOL (Jullo et al., 2007) to measure the enclosed total mass—as opposed to the simple evaluation of Equations (1), (2), and (3). For the first 35 clusters that were analyzed in this work, we generated single-halo lens models by taking the three coordinates in each lensed arc used to derive the Einstein radius in Section 2.1 as a multiple-image family with a single pseudo-isothermal elliptical mass distribution (PIEMD, Kassiola & Kovner 1993) locked to the center of the BCG. Mass estimates were then obtained from the resulting best-fit lens models according to the methodology described in Remolina González et al. (2021a).
Figure 5: A comparison of the single-halo LENSTOOL-generated mass measurements ($M_{SHM}$) and the parametrically generated mass measurements ($M_{\Sigma}$) for the first 35 clusters analyzed in this work. The black line represents an ideal 1-1 correspondence, and the blue lines represent the uncertainty on the measured correspondence.
As expected from Remolina González et al. (2021b), these differing methodologies show similar results. The two mass measurements for the 35 clusters studied here are correlated with a slope of $0.960\pm 0.036$ and intercept of $0.492\pm 0.476$ dex as shown in Figure 5. In other words, we find no reason to suggest a preference between one method over the other, and we abandoned the LENSTOOL approach in favor of the much faster parametric approach throughout the rest of the work.
In theory, lens modelling has the potential to be the most accurate mass-constraining methodology. However, this requires high-spatial-resolution imaging data and ample time. The process of iteratively refining just one lens model may take upwards of an hour for even a single relatively simple system—which is untenable when the number of candidate systems reaches into the thousands. Moreover, the accuracy of measuring mass with lens modelling is sensitive to the choice of defining multiple-image families along the lensed source arcs in a way that simple circle fitting is not. These essential positional constraints are often unresolved even with the highest-quality ground-based imaging data. Conversely, the positional constraints that inform the Einstein radius ultimately used by Equations (1), (2), and (3) need not be extremely accurate with respect to exactly constraining image family positions. As long as the positional constraints are reasonably placed within the arcs, something which is resolved by ground-based imaging data, it is accurate to within the systematic uncertainty.
For more information on lens-modelling software, related techniques, and applications of strong lensing, we refer the reader to the following non-exhaustive list of papers for further review: Elíasdóttir et al. (2007), Oguri (2010), Lefor et al. (2013), Meneghetti et al. (2017), Birrer & Amara (2018), Schäfer et al. (2020), Sharon et al. (2022), and Napier et al. (2023a).
@@ 6 Summary and Conclusions
This paper presents measurements of mass and light in 177 strong-lensing systems at cluster and group mass scales. By directly obtaining the Einstein radius and BCG redshift for each system, we compute the Einstein-radius-enclosed total mass parametrically. We use the enclosed $g$, $r$, and $z$-band photometry with Prospector to parameterize the enclosed luminosity and stellar mass, and we obtain a planar relationship between the BCG redshift, enclosed total mass, and enclosed luminosity of the form $Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$. We find that the enclosed total mass and stellar mass for the entire sample of strong-lensing systems considered in this work are correlated with a logarithmic slope, normalization, and intrinsic stellar-mass scatter of $0.443\pm 0.035$, $5.958\pm 0.454$ dex, and $0.173\pm 0.013$ dex respectively. We also find that the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope, normalization, and intrinsic stellar-to-total-mass-fraction scatter of $-0.563\pm 0.035$, $6.034\pm 0.457$ dex, and $0.161\pm 0.013$ dex respectively. Enormous volumes of imaging data from new space-based and ground-based surveys at deeper and sharper photometric limits are on the horizon. Probing these upcoming datasets with both visual and machine-learning driven searches for new strong lenses at a variety of scales is inevitable. The correlations shown in this work should have utility in filtering rankings of possible strong-lensing systems as well as offering another measure of cluster cores that can be compared to cosmological simulations.